# LLMs may Dominate Information Access: Neural Retrievers are Biased Towards LLM-Generated Texts

### v2

@@ Abstract.
Recently, the emergence of large language models (LLMs) has revolutionized the paradigm of information retrieval (IR) applications, especially in web search. With their remarkable capabilities in generating human-like texts, LLMs have created enormous texts on the Internet. As a result, IR systems in the LLMs era are facing a new challenge: the indexed documents now are not only written by human beings but also automatically generated by the LLMs. How these LLM-generated documents influence the IR systems is a pressing and still unexplored question. In this work, we conduct a quantitative evaluation of different IR models in scenarios where both human-written and LLM-generated texts are involved. Surprisingly, our findings indicate that neural retrieval models tend to rank LLM-generated documents higher. We refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. Moreover, we discover that this bias is not confined to the first-stage neural retrievers, but extends to the second-stage neural re-rankers. Then, we provide an in-depth analysis from the perspective of text compression and observe that neural models can better understand the semantic information of LLM-generated text, which is further substantiated by our theoretical analysis. To mitigate the source bias, we also propose a plug-and-play debiased constraint for the optimization objective, and experimental results show the effectiveness. Finally, we discuss the potential severe concerns stemming from the observed source bias and hope our findings can serve as a critical wake-up call to the IR community and beyond. To facilitate future explorations of IR in the LLM era, the constructed two new benchmarks and codes will later be available at https://github.com/KID-22/LLM4IR-Bias.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: XXXX; August 03–05, 2023; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Meanwhile, information retrieval (IR) systems have become the keystone of how individuals navigate and access information within the vast expanse of the Internet (Singhal et al., 2001; Manning, 2009). The core function of an IR system is to efficiently retrieve relevant documents from the target corpus in response to users’ queries (Liu et al., 2009; Li, 2022; Xu and Li, 2007). As illustrated in Figure 1, in the era preceding the widespread emergence of LLMs, IR systems primarily revolved around the retrieval of relevant documents solely from the human-written corpus. However, with the rise of LLMs, IR systems are confronted with a profound paradigm shift. That is, as LLMs have significantly driven the proliferation of AIGC, the corpus of IR systems now gradually includes both human-written and LLM-generated texts. Consequently, this paradigm shift raises a fundamental research question: What is the impact of the proliferation of generated content on IR systems? In this LLM era, we aim to explore whether existing retrieval models tend to prioritize LLM-generated text over human-written text, even when both convey similar semantic information. If this holds, LLMs may dominate information access, particularly as their generated content is rapidly growing on the Internet (Hanley and Durumeric, 2023).
@@
To approach the fundamental research question, we decompose it into three specific research questions. The first question is RQ1: How to construct a simulation environment to evaluate IR models in the LLM era? As the LLM-generated content is currently unidentifiable on the Internet, there are no public retrieval benchmarks where the corpus contains both human-written text and LLM-generated text. To construct a more realistic evaluation environment, we introduce a natural approach to synthesize LLM-generated data without incurring significant costs for human annotation of relevant labels. Specifically, we leverage the original human-written texts as the instruction conditions to prompt LLMs to generate rewritten text copies while preserving the same semantic meaning. In this way, we can confidently assign the relevant labels for LLM-generated data. Extensive empirical analysis validates the quality of our constructed environment, as it can accurately reflect real-world situations of IR in the LLM era. As a result, we introduce two new benchmarks, SciFact+AIGC and NQ320K+AIGC, tailored for IR research in the LLM era.
@@
Based on the constructed environment, we can explore further RQ2: Are and where are retrieval models biased towards LLM-generated texts? We conduct systematical experiments with various representative and state-of-the-art retrieval models, ranging from traditional lexical models to modern neural retrieval models (Zhao et al., 2022; Yates et al., 2021; Guo et al., 2022). Surprisingly, our findings indicate that neural retrieval models are biased towards LLM-generated texts, i.e., rank LLM-generated texts in a higher position. We refer to this as source bias, as the neural retrievers favor texts from specific sources (i.e., LLM-generated). Our further experiments indicate that the source bias not only extends to the second-stage neural re-rankers but also manifests in a more severe manner. These findings further corroborate the prevalence of source bias in neural ranking models that rely on pretrained language models (PLMs) (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022). In other words, these neural ranking models appear to adopt some inherent semantic shortcuts to learn the semantic representation of documents generated by LLM.
@@
Then, what we are curious about is RQ3: Why are neural retrieval models biased towards LLM-generated texts? Inspired by the recent studies that posit LLMs are lossless compressors (Delétang et al., 2023), we aim to analyze the cause of source bias from the viewpoint of text compression. We first examine the singular value (Klema and Laub, 1980) of different corpora and uncover that LLM-generated text typically maintains more focused semantics with minimal noise, rendering it better understood for precise semantic matching. Furthermore, to better examine whether the neural retrieval model based on PLMs can better understand the LLM-generated texts, we conduct an in-depth analysis for perplexity. Our experimental and theoretical analysis suggest that LLM-generated text consistently yields lower perplexity scores, signifying greater comprehensibility and higher confidence from the perspective of PLMs. These observations confirm that LLM-generated text is more readily understandable to neural retrieval models, thereby resulting in source bias.
@@
As our analysis has revealed source bias, we aim to further study the RQ4: How can we mitigate the source bias in neural retrieval models? Specifically, we propose a simple but effective debiased constraint to penalize biased samples during optimization. In this way, we can force the retrieval models to focus on modeling semantic relevance rather than the inherent shortcut between neural retrievers and LLM-generated content, leading to the mitigation of source bias. This debiased constraint can be plugged and played into the ranking optimization objective of various neural retrieval models. Besides, our approach offers the capability to control the degree of bias removal, providing flexibility in determining whether the two sources of content should be treated equally or biased toward human-written content according to specific requirements and environmental considerations.
@@
Finally, we discuss the potential emerging concerns stemming from source bias. First, the presence of source bias poses a significant risk of gradually rendering human-written content less accessible, potentially causing a disruption in the content ecosystem. This concern is further compounded by the growing prevalence of LLM-generated text on the Internet(Hanley and Durumeric, 2023; Bengio et al., 2023). Second, there is the risk that source bias may amplify the spread of misinformation, especially considering the potential of LLMs to generate deceptive content, whether intentionally or not (Chen and Shu, 2023; Su et al., 2023). Third, source bias may be maliciously exploited to manipulate existing retrieval and ranking algorithms, creating a precarious vulnerability that could be weaponized by malicious actors. In light of these pressing issues, we hope that our findings serve as a resounding wake-up call to all stakeholders involved in IR systems and beyond.
@@
(1) This is the first work that introduces the paradigm of IR systems in the LLM era, where the retrieval corpus consists of both human-written texts and LLM-generated texts. This scenario is a more realistic setting, considering the growing prosperity of AIGC.
(2) We uncover the source bias in the LLM era, a new inherent bias in neural retrieval models, where both the neural retrievers and re-rankers rank the LLM-generated text in a higher position than human-written text with the almost same semantic relevance.
(3) We provide an in-depth analysis and theoretical insights of source bias from the viewpoint of text compression. The analysis indicates that LLM-generated text maintains more focused semantics with minimal noise and is more readily comprehensible for neural retrieval models.
(4) We propose a debiased constraint to penalize the bias of neural retrieval models during the optimization and experimental results demonstrate the effectiveness in mitigating source bias.
(5) We also provide two new benchmarks, SciFact+AIGC and NQ320K+AIGC, which contain both high-quality human-written and various LLM-generated corpus and corresponding relevant labels. We believe these two benchmarks can serve as valuable resources for facilitating future research of IR in the LLM era.
@@ 2. RQ1: Environment Construction
With the increasing usage of LLMs in generating texts (e.g., paraphrasing or rewriting texts with LLMs), the corpus of IR systems includes both human-written and LLM-generated texts nowadays. In this section, we introduce a novel and natural framework for the quantitative evaluation of retrieval models in the LLM era, i.e., in scenarios where the corpus includes both human-written and LLM-generated texts, as shown in Figure 2. To better align with the real-world scenarios in the LLM era, the simulated environments should meet the following two essential criteria. Firstly, it is imperative to distinguish between human-written texts and LLM-generated texts within the corpus. Secondly, we need access to pertinent labels for LLM-generated data in response to queries. To ensure the most effective evaluation of retrieval model preferences, every piece of human-written text should have a corresponding LLM-generated counterpart with consistent content.
2.1. Preliminary
2.1.1. Notation
Formally, in the Pre-LLM era, given a query text $q\in\mathcal{Q}$ where $\mathcal{Q}$ is the set of all queries, the traditional IR system aims to retrieve a list of top-$K$ relevant documents $\{d^{(1)},d^{(2)},\ldots,d^{(K)}\}$ from a corpus $\mathcal{C}^{H}=\{d_{1}^{H},d_{2}^{H},\ldots d_{N}^{H}\}$ which consists of $N$ human-written documents. However, in the era of LLMs, there is also LLM-generated text in the corpus. To simulate and evaluate the IR models in the LLM era, we also create an additional corpus $\mathcal{C}^{G}=\{d_{1}^{G},d_{2}^{G},\ldots,d_{N}^{G}\}$ where each document is generated by a LLM, e.g., $d_{1}^{G}$ can be created by ChatGPT by constructing a prompt that asks ChatGPT to rewrite $d_{1}^{H}$ while preserving its original meaning. Consequently, given a user query $q$, the primary objective of a retriever in the LLM era is to return the top-$K$ relevant documents from the combined corpus $\mathcal{C}=\mathcal{C}^{H}\bigcup\mathcal{C}^{G}$.
2.1.2. Retrieval Models
IR models have witnessed a significant evolution, shifting from traditional lexical approaches primarily reliant on term matching to the recent neural IR models founded on semantic matching (Zhao et al., 2022; Guo et al., 2022; Yates et al., 2021; Gonzalo et al., 2014). In this study, we select the following representative retrieval models to conduct a thorough exploration of their performance in scenarios encompassing both human-written and LLM-generated texts:
Lexical Retrieval Models are based on the exact term matching between high-dimensional sparse vectors of query and document. We choose the widely used two approaches TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009) in our experiments.
Neural Retrieval Models utilize neural networks to learn semantic relationships between query and document to improve the lexical gap (Li et al., 2014; Zhao et al., 2022; Yates et al., 2021; Guo et al., 2020). We select the following four representative neural retrieval models: ANCE (Xiong et al., 2020) is a dense retrieval model that selects hard training negatives globally from the entire corpus, using an asynchronously updated ANN index. BERM (Xu et al., 2023) is also a dense retrieval model that focuses on learning the domain-invariant features in information retrieval tasks. TAS-B (Hofstätter et al., 2021) is a bi-encoder trained with a balanced margin sampling technique that samples queries out of a cluster per batch, ensuring efficiency in the sampling process. Contriever (Izacard et al., 2021) uses momentum contrastive learning to train dense retrievers from text paragraphs and obtain state-of-the-art zero-shot retrieval performance.
@@ 2.1.2. Retrieval Models
Dataset # Test Queries # Avg. Query Length Human-Written Corpus Llama2-Generated/ChatGPT-Generated Corpus
@@
SciFact+AIGC 300 12.38 5,183 201.81 1.1 5,183/5,183 192.66/203.57 1.1/1.1
NQ320K+AIGC 7,830 9.24 109,739 199.79 1.0 109,739/- 174.49/- 1.0/-
@@ 2.2. Constructing IR Datasets in the LLM Era
To better align with real-world IR scenarios in the LLM era, we first prompt LLMs to rewrite human-written corpus to build two new standard retrieval datasets: SciFact+AIGC and NQ320K+AIGC. These two new datasets can serve as valuable resources to facilitate future research of IR in the LLM era.
(a) Jaccard Similarity
(b) Overlap
(a) SciFact+AIGC
(b) NQ320K+AIGC
2.2.1. Human-Written Corpus
We first choose two widely used retrieval datasets written by humans in the Pre-LLM era as the seed data: SciFact and NQ320K. SciFact [^1] https://allenai.org/data/scifact (Wadden et al., 2020) dataset aims to retrieve evidence from the research literature containing scientific paper abstracts for fact-checking. NQ320K [^2] https://ai.google.com/research/NaturalQuestions (Kwiatkowski et al., 2019) is based on the Natural Questions (NQ) dataset from Google, where the documents are gathered from Wikipedia pages, and the queries are natural language questions. Following the practice in BEIR benchmark (Thakur et al., 2021), we process these two datasets in a standard format: corpus $\mathcal{C}^{H}$, queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$, where $M$ is the number of labeled query-document pairs in the corpus.
2.2.2. LLM-Generated Corpus
Constructing an IR dataset in the LLM era typically involves two steps: collecting the LLM-generated corpus $\mathcal{C}^{G}$ and employing human evaluators from crowdsourcing platforms to annotate relevance labels for each query-document pair. Given the significant cost associated with human annotation, we adopt an alternative natural approach. Instead, we repurpose the original human-written corpus as our seed data source and directly instruct LLMs to rewrite each given text from the human-written corpus. As the written text generated by LLM carries almost the same semantic information as the original human-written text, we can assign the same relevance labels to new query-document pairs as those assigned to the original query-document pairs.
Our instruction is straightforward: “Please rewrite the following text: {{human-written text}}”, as illustrated in the left part of Figure 2. This straightforward instruction enables LLMs to generate text without too many constraints while maintaining semantic equivalence to the original human-written text. Specifically, we choose Llama2 (Touvron et al., 2023) and ChatGPT to rewrite each seed human-written corpus, as Llama2 and ChatGPT are both the most widely-used and nearly the state-of-the-art open-sourced and closed-source LLM, respectively. We only generate texts with ChatGPT corresponding to the texts in SciFact dataset, mainly due to the significant cost involved in processing the larger NQ320K dataset. If not specified, the temperature of the used LLM for generation corpus is set as 0.2 by default for the following experiments.
As a result, we can obtain two corresponding LLM-generated corpora with SciFact and NQ320K as seed data. After that, we extend the original labels of query and human-written text $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$ to get the corresponding label of LLM-generated text $\mathcal{R}^{G}=\{(q_{m},d^{G}_{m},r_{m})\}^{M}_{m=1}$. We will validate the quality of the datasets in the following section. Combining each original human-written corpus $\mathcal{C}^{H}$ with its corresponding LLM-generated corpus $\mathcal{C}^{G}$, original queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}\bigcup\mathcal{R}^{G}$, we can create two new datasets, denoted as SciFact+AIGC and NQ320K+AIGC.  Table 1 summarize the statistics of the proposed two datasets.

### v3

# Neural Retrievers are Biased Towards LLM-Generated Content

@@ Abstract.
Recently, the emergence of large language models (LLMs) has revolutionized the paradigm of information retrieval (IR) applications, especially in web search, by generating vast amounts of human-like texts on the Internet. As a result, IR systems in the LLM era are facing a new challenge: the indexed documents are now not only written by human beings but also automatically generated by the LLMs. How these LLM-generated documents influence the IR systems is a pressing and still unexplored question. In this work, we conduct a quantitative evaluation of IR models in scenarios where both human-written and LLM-generated texts are involved. Surprisingly, our findings indicate that neural retrieval models tend to rank LLM-generated documents higher. We refer to this category of biases in neural retrievers towards the LLM-generated content as the source bias. Moreover, we discover that this bias is not confined to the first-stage neural retrievers, but extends to the second-stage neural re-rankers. Then, in-depth analyses from the perspective of text compression indicate that LLM-generated texts exhibit more focused semantics with less noise, making it easier for neural retrieval models to semantic match. To mitigate the source bias, we also propose a plug-and-play debiased constraint for the optimization objective, and experimental results show its effectiveness. Finally, we discuss the potential severe concerns stemming from the observed source bias and hope our findings can serve as a critical wake-up call to the IR community and beyond. To facilitate future explorations of IR in the LLM era, the constructed two new benchmarks are available at https://github.com/KID-22/Source-Bias.
Source Bias, Information Retrieval, LLM-Generated Texts, Artificial Intelligence Generated Content
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671882isbn: 979-8-4007-0490-1/24/08ccs: Information systems Information retrieval
@@ 1. Introduction
Meanwhile, information retrieval (IR) systems have become indispensable for navigating and accessing the Internet’s vast information landscape (Singhal et al., 2001; Manning, 2009). As illustrated in Figure 1, in the era preceding the widespread emergence of LLMs, IR systems focused on retrieving documents solely from the human-written corpus in response to users’ queries (Liu et al., 2009; Li, 2022; Xu and Li, 2007). However, the proliferation of AIGC driven by LLMs has expanded the corpus of IR systems to include both human-written and LLM-generated texts. Consequently, this paradigm shift raises a fundamental research question: What is the impact of the proliferation of generated content on IR systems? We aim to explore whether existing retrieval models tend to prioritize LLM-generated text over human-written text, even when both convey similar semantic information. If this holds, LLMs may dominate information access, particularly as their generated content is rapidly growing on the Internet (Hanley and Durumeric, 2023).
@@
To approach the fundamental research question, we decompose it into four specific research questions. The first question is RQ1: How to construct an environment to evaluate IR models in the LLM era? Given the lack of public retrieval benchmarks encompassing both human-written and LLM-generated texts, we propose an innovative and practical method to create such a realistic evaluation environment without the need of costly human annotation. Specifically, we leverage the original human-written texts as the instruction conditions to prompt LLMs to generate rewritten text copies while preserving the same semantic meaning. In this way, we can confidently assign the same relevancy labels to LLM-generated data according to the original labels. Extensive empirical analysis validates the quality of our constructed environment, demonstrating its effectiveness in mirroring real-world IR scenarios in the LLM era. As a result, we introduce two new benchmarks, SciFact+AIGC and NQ320K+AIGC, tailored for IR research in the LLM era.
@@
With the constructed environment, we further explore RQ2: Are retrieval models biased towards LLM-generated content? We conduct comprehensive experiments with various representative retrieval models, ranging from traditional lexical models to modern neural models based on pretrained language models (PLMs) (Guo et al., 2020; Zhao et al., 2023a; Yates et al., 2021; Guo et al., 2022). Surprisingly, we uncover that neural retrievers are biased towards LLM-generated texts, i.e., tend to rank LLM-generated texts in higher positions. We refer to this as source bias, as the neural retrievers favor content from specific sources (i.e., LLM-generated content). Further experiments indicate that the source bias not only extends to the second-stage neural re-rankers from the first-stage retrieval but also manifests more severely. These findings corroborate the prevalence of source bias in neural retrieval models.
@@
Then, what we are curious about is RQ3: Why are neural retrieval models biased towards LLM-generated texts? Inspired by the recent studies positing LLMs as lossless compressors (Delétang et al., 2023), we analyze the cause of source bias from the viewpoint of text compression. Our analysis of singular values (Klema and Laub, 1980) in different corpora reveals that LLM-generated texts exhibit more focused semantics with minimal noise, enhancing their suitability for semantic matching. Furthermore, our in-depth perplexity analysis shows that LLM-generated texts consistently achieve lower perplexity scores, which indicates a higher degree of comprehensibility and confidence from the PLM’s perspective. These observations collectively suggest that LLM-generated texts are more readily understandable to PLM-based neural retrievers, thereby resulting in source bias.
@@
Finally, we try to answer RQ4: How to mitigate source bias in neural retrieval models? To tackle this, we propose an intuitive yet effective debiased constraint. This constraint is designed to penalize biased samples during the optimization process, thereby shifting the focus of retrieval models from exploiting inherent shortcuts to emphasizing semantic relevance. Besides, our debiased constraint is model-agnostic and can be plugged and played to the ranking optimization objectives of various neural retrieval models. Furthermore, it offers the capability to control the degree of bias removal, offering the flexibility to balance the treatment between the two sources of content based on specific requirements and environmental considerations.
@@
Last but not least, we discuss the potential emerging concerns stemming from source bias, highlighting the risk of human-written content being gradually inaccessible, especially due to the rapidly increasing LLM-generated content on the Internet (Hanley and Durumeric, 2023; Bengio et al., 2023). Furthermore, source bias could be maliciously exploited to manipulate algorithms and potentially amplify the spread of misinformation, posing a threat to online security. In light of these pressing issues, we hope that our findings serve as a resounding wake-up call to all stakeholders involved in IR systems and beyond.
@@
(1) We introduce a more realistic paradigm of IR systems considering the growing prosperity of AIGC, where the retrieval corpus consists of both human-written and LLM-generated texts. We then uncover a new inherent bias in both neural retrieves and re-rankers preferring LLM-generated content, termed as source bias.
(2) We provide an in-depth analysis and insights of source bias from a text compression perspective, which indicates that LLM-generated texts maintain more focused semantics with minimal noise and are more readily comprehensible for neural retrievers.
(3) We propose a debiased constraint to penalize the biased samples during optimization, and experimental results demonstrate its effectiveness in mitigating source bias in different degrees.
(4) We also provide two new benchmarks, SciFact+AIGC and NQ320K+AIGC, which contain both high-quality human-written and various LLM-generated corpus and corresponding relevancy labels. We believe these two benchmarks can serve as valuable resources for facilitating future research of IR in the LLM era.
@@ 2. RQ1: Environment Construction
With the increasing usage of LLMs in generating texts (e.g., paraphrasing or rewriting), the corpus of IR systems includes both human-written and LLM-generated texts nowadays. Constructing an IR dataset in the LLM era typically involves two steps: collecting both human-written and LLM-generated corpora and then employing human evaluators to annotate relevancy labels for each query-document pair. Given that LLM-generated content is currently unidentifiable (Sadasivan et al., 2023) and the significant cost of human annotation, we introduce a natural and practical framework for quantitatively evaluating retrieval models in the LLM era, as shown in Figure 2.
To better align with real-world scenarios, the evaluation environments should meet the following three essential criteria. Firstly, it is imperative to distinguish between human-written and LLM-generated texts within the corpus. Secondly, we need access to relevancy labels for LLM-generated data in response to queries. Thirdly, each human-written text should better have a corresponding LLM-generated counterpart with the same semantics, ensuring the most effective and fair evaluation.
@@
Dataset # Test Queries # Avg. Query Length Human-Written Corpus Llama2-Generated Corpus ; ChatGPT-Generated Corpus
@@
SciFact+AIGC 300 12.38 5,183 201.81 1.1 5,183 ; 5,183 192.66 ; 203.57 1.1 ; 1.1
NQ320K+AIGC 7,830 9.24 109,739 199.79 1.0 109,739 ; – 174.49 ; – 1.0 ; –
2.1. Notation
Formally, in the Pre-LLM era, given a query $q\in\mathcal{Q}$ where $\mathcal{Q}$ is the set of all queries, the traditional IR system aims to retrieve a list of top-$K$ relevant documents $\{d^{(1)},d^{(2)},\ldots,d^{(K)}\}$ from a corpus $\mathcal{C}^{H}=\{d_{1}^{H},d_{2}^{H},\ldots d_{N}^{H}\}$ which consists of $N$ human-written documents. However, in the era of LLMs, there is also LLM-generated text in the corpus. To evaluate the IR models in the LLM era, we also create an additional LLM-generated corpus $\mathcal{C}^{G}=\{d_{1}^{G},d_{2}^{G},\ldots,d_{N}^{G}\}$ where each document is generated by a LLM, e.g., $d_{1}^{G}$ can be created by prompting ChatGPT to rewrite $d_{1}^{H}$ while preserving its original semantics information. Consequently, given a query $q$, the objective of a retriever in the LLM era is to return the top-$K$ relevant documents from the mixed corpus $\mathcal{C}=\mathcal{C}^{H}\bigcup\mathcal{C}^{G}$.
@@ 2.2. Constructing IR Datasets in the LLM Era
In this section, we prompt LLMs to rewrite human-written corpus to build two new standard retrieval datasets: SciFact+AIGC and NQ320K+AIGC. These two new datasets can serve as valuable resources to facilitate future research of IR in the LLM era.
@@
2.2.1. Human-Written Corpus
We first choose two widely used retrieval datasets written by humans in the Pre-LLM era as the seed data: SciFact and NQ320K. SciFact [^1] https://allenai.org/data/scifact (Wadden et al., 2020) dataset aims to retrieve evidence from the research literature containing scientific paper abstracts for fact-checking. NQ320K [^2] https://ai.google.com/research/NaturalQuestions (Kwiatkowski et al., 2019) is based on the Natural Questions (NQ) dataset from Google, where the documents are gathered from Wikipedia pages, and the queries are natural language questions. Following the practice in BEIR benchmark (Thakur et al., 2021), we process these two datasets in a standard format: corpus $\mathcal{C}^{H}$, queries $\mathcal{Q}$, and relevancy labels $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$, where $M$ is the number of labeled query-document pairs in the dataset.
2.2.2. LLM-Generated Corpus
For the LLM-generated corpus, we repurpose the original human-written corpus as our seed data and instruct LLMs to rewrite each given text from the human-written corpus. As the written text generated by LLM carries almost the same semantic information as the original human-written text, we can assign the same relevancy labels to new <query, LLM-generated document> pairs as those assigned to the original labeled <query, human-written document> pairs.
Our instruction is straightforward: “Please rewrite the following text: {{human-written text}}”, as illustrated in the left part of Figure 2. This straightforward instruction enables LLMs to generate text without too many constraints while maintaining semantic equivalence to the original human-written text. Specifically, we choose Llama2 (Touvron et al., 2023) and ChatGPT to rewrite each seed human-written corpus, as Llama2 and ChatGPT are both the most widely-used and nearly the state-of-the-art open-sourced and closed-source LLM, respectively. We only generate texts with ChatGPT corresponding to the texts in SciFact dataset, mainly due to the significant cost involved in processing the larger NQ320K dataset.
For the LLM-generated corpus, we conduct post-processing to remove unrelated parts of the original response from LLM like “Sure, here’s a possible rewrite of the text:”. As a result, we can obtain two corresponding LLM-generated corpora with SciFact and NQ320K as seed data. After that, we extend the original labels of query and human-written text $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$ to get the corresponding label of LLM-generated text $\mathcal{R}^{G}=\{(q_{m},d^{G}_{m},r_{m})\}^{M}_{m=1}$. We will validate the quality of the datasets in the following section. Combining each original human-written corpus $\mathcal{C}^{H}$ with its corresponding LLM-generated corpus $\mathcal{C}^{G}$, original queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}\bigcup\mathcal{R}^{G}$, we can create two new datasets, denoted as SciFact+AIGC and NQ320K+AIGC.  Table 1 summarizes the statistics of the proposed two datasets.
(a) Jaccard Similarity
(b) Overlap